OpenStudy (caroline1218):

What is (0,-2) and (3,2) in standard form, point slope form, and slope-intercept form?

OpenStudy (caroline1218):

@pooja195 @imqwerty @jhonyy9 @Mehek14 @nincompoop

OpenStudy (caroline1218):

I know the formulas

OpenStudy (caroline1218):

I'm not sure how to get the standard form

OpenStudy (caroline1218):

Point Slope Form is y = 4x/3 - 2. Slope-intercept Form is y = 4/3x -2.

OpenStudy (caroline1218):

So now how do I get standard form @Mehek14 ?

OpenStudy (caroline1218):

And were my point slope form and slope-intercept form correct?

OpenStudy (caroline1218):

Ok, but point slope form is wrong?

OpenStudy (caroline1218):

Is that the correct answer for point-slope form?

OpenStudy (caroline1218):

Ok, so then what is the standard form?

OpenStudy (caroline1218):

Ok

OpenStudy (caroline1218):

Then what?

OpenStudy (caroline1218):

so y divided by 4/3 = x - 2 ?

OpenStudy (caroline1218):

I'm confused

OpenStudy (caroline1218):

So 1 1/3 + -1y = 2 ?

OpenStudy (caroline1218):

@Mehek14

OpenStudy (serenity74):

yes

OpenStudy (caroline1218):

Yes that is the correct standard form?

OpenStudy (serenity74):

NO that's wrong actually 1*4/3 = ?

OpenStudy (caroline1218):

Oh, so just 4/3 + -1y = 2 ?

OpenStudy (serenity74):

no 4/3x - y =2

OpenStudy (caroline1218):

Ok, thank you :)

OpenStudy (serenity74):

:)